OpenStudy (anonymous):

4. Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 18x + 6y = 78 12x + 54y = -48

OpenStudy (anonymous):

i got 30x+60y=22

OpenStudy (anonymous):

Give me a few min. - I'm working on it.......

OpenStudy (amistre64):

(18x + 6y = 78)/6 3x + y = 13 (12x + 54y = -48)/6 2x +9y = -8 3x + y = 13 2x +9y = -8

OpenStudy (amistre64):

those are equavialent and easier equations to work with

OpenStudy (amistre64):

it asking you what value for x and y make the system consistent from the first eq we can say that: y = 13-3x and substitute that "almost" value into the second eq and solve that for x

OpenStudy (amistre64):

2x +9y = -8 2x +9(13-3x) = -8 2x +9(13) -27x = -8 2x -27x = -8 -9(13) -25x = -8 -9(13) x = \(\cfrac{-8-9(13)}{-25}\) = \(\cfrac{125}{25}\)=5

OpenStudy (amistre64):

since x = 5, if i did it right, we can solve for y now y = 13 - 3x y = 13 - 3(5) y = 13 - 15 = -2

OpenStudy (amistre64):

its always good to dbl chk the results since head math is rife with error

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

18x + 6y = 78 5 -2 90 - 12 = 78 .... good 12x + 54y = -48 5 -2 60 - 108 = -48 .... good

OpenStudy (anonymous):

@amistre64 - Could you please check the answer I got on my last question - I am still confused with the answer I got, thanks :)